# Quantum adiabatic optimization with Rydberg arrays: localization phenomena and encoding strategies

### v1

@@ Abstract
We study the quantum dynamics of the encoding scheme proposed in [Nguyen et al., PRX Quantum 4, 010316 (2023)], which encodes optimization problems on graphs with arbitrary connectivity into Rydberg atom arrays. Here, a graph vertex is represented by a wire of atoms, and the (crossing) crossing-with-edge gadget is placed at the intersection of two wires to (de)couple their degrees of freedom and reproduce the graph connectivity. We consider the fundamental geometry of two vertex-wires intersecting via a single gadget and look at minimum gap scaling with system size along adiabatic protocols. We find that both polynomial and exponential scaling are possible and, by means of perturbation theory, we relate the exponential closing of the minimum gap to an unfavorable localization of the ground-state wavefunction. Then, on the QuEra Aquila neutral atom machine, we observe such localization and its effect on the success probability of finding the correct solution to the encoded optimization problem. Finally, we propose possible strategies to avoid this quantum bottleneck, leading to an exponential improvement in the adiabatic performance.
@@ I Introduction
The past few decades have seen a growing interest in harnessing the dynamics of many-body quantum systems to explore potential quantum advantage in solving combinatorial optimization problems [1, 2, 3, 4]. For problems that can be reformulated in terms of a classical spin Hamiltonian with ground state encoding their solution, quantum adiabatic optimization stands out as one of the most promising quantum algorithms for pursuing such advantage [5, 6, 7, 8, 9, 10, 11]. Here, the Hamiltonian of the system is adiabatically evolved from one with a known and easy-to-prepare ground state to the target one. In this way, the system follows the instantaneous ground state of the time-dependent Hamiltonian throughout the evolution and finally reaches the solution to the optimization problem. Among the various platforms, Rydberg atom arrays have recently generated interest for implementing such optimization [12, 4, 13, 14, 15, 16, 17]. Due to the blockade mechanism, they naturally encode the paradigmatic maximum independent set problem on a specific class of graphs, so-called unit-disk graphs [12], and experiments have already shown promising results [4]. Developing strategies to go beyond the native hardware restriction of unit-disk graphs is of central importance for broadening the class of graphs and problems that can be investigated using this platform  [18, 19, 20, 21, 22, 23, 24].
@@
To address this challenge, Ref. [19] proposed an encoding strategy that maps an optimization problem on an arbitrarily connected graph into an equivalent maximum weighted independent set (MWIS) problem on a unit-disk graph, which is directly implementable with Rydberg atom arrays (see Fig. 1). Here, the binary degree of freedom associated with each graph vertex is encoded in the $\mathbb{Z}_{2}$ ordered states of a vertex-wire of Rydberg atoms. The absence (presence) of an edge between a pair of vertices is reproduced by placing a crossing (crossing-with-edge) gadget at the intersection between the corresponding vertex-wires. So far, this strategy has been mainly investigated on the classical level, and the quantum adiabatic performance, in particular for large systems, remains an outstanding open question.
@@
In this work, we investigate the adiabatic performance of the main building blocks of this encoding scheme, consisting of two vertex-wires intersecting either via the crossing or the crossing-with-edge (CWE) gadget, as depicted in Fig. 1(c,d). When the two vertex-wires intersect via the crossing-with-edge gadget, we find that an exponential closing of the minimum gap can occur for certain geometrical configurations, leading to exponentially long algorithm runtimes for large systems. We show that this adversarial gap closing originates from unfavorable localization of the ground-state wavefunction along the adiabatic protocol. Guided by these insights, we propose alternative strategies to overcome such localization and ensure a polynomial closing of the minimum gap, thereby exponentially improving the adiabatic performance of the encoding scheme. In particular, we show how localization can be avoided by (i) systematically varying the length of the intersecting vertex-wires, or (ii) modifying the structure of the crossing-with-edge gadget.
@@
This work is organized as follows. In Sec. II, we revisit the encoding of the MWIS problem on arbitrarily connected graphs into Rydberg atom arrays, as proposed in Ref. [19]. After detailing its building blocks (namely, the crossing and the CWE setup), we introduce the two adiabatic protocols that we consider in this work, referred to as standard and logical protocols. In Sec. III, we employ the density matrix renormalization group (DMRG) method to quantify the adiabatic performance of the two setups along the standard protocol. In Sec. IV, we develop an effective theory to characterize their performance along the logical protocol and explain the origin of the wavefunction localization. In Sec. V, using the QuEra Aquila neutral atom machine [25], we observe such localization phenomenon and its effect on encoded MWIS problems. Finally, in Sec. VI, we present the strategies to overcome the localization of the ground-state wavefunction and achieve an exponential improvement in the adiabatic performance.
@@ II Rydberg encoding of maximum weighted independent set problems
Figure 1: Encoding scheme of the maximum weighted independent set problem on arbitrarily connected graphs into Rydberg atom arrays [19]. (a) Example for a non-unit-disk graph of five vertices $G=(V,E,\delta)$. (b) A vertex-wire is a one-dimensional array of nearest-neighbor-blockaded atoms. (left) The atoms are labeled from $1$ to $N$, with $N$ even. The detuning in the bulk is denoted by $\Delta$; the detuning on the initial and final atom by $\Delta-\epsilon_{i}$ and $\Delta-\epsilon_{f}$. The weight of a graph vertex $\delta$ is mapped to the boundary detuning difference: $\delta=\epsilon_{f}-\epsilon_{i}$. (right) The logical states encoding the binary degrees of freedom of a graph vertex are the $\mathbb{Z}_{2}$ ordered states of the vertex-wire. Here, the atoms in the ground (Rydberg) state are depicted in white (black). The (c) crossing and (d) crossing-with-edge (CWE) setups consist of two vertex-wires intersecting via the crossing gadget (the four atoms in gray) and the CWE gadget (the three atoms in gray, labeled as in the inset). (left) The atoms along the horizontal ($w=x$) and vertical ($w=y$) vertex-wire are labeled from $1$ to $N_{w}$; the position of the last atom before the gadget is denoted by $m_{w}$. $N_{w}$ and $m_{w}$ are chosen to be even. The bulk detuning is denoted by $\Delta$; the detuning on the initial and the final atom of the $w$ vertex-wire by $\Delta-\epsilon_{i,w}$ and $\Delta-\epsilon_{f,w}$; the detuning on the gadget atoms by $\Delta_{g}$. (right) The logical states of the crossing setup are the four possible combinations of logical states of the vertex-wires. In the CWE setup, the logical state $|11\rangle_{l}$ is forbidden due to the blockade mechanism.
@@ II.2 MWIS encoding for arbitrarily connected graphs
A quadratic overhead in the atom number makes it possible to go beyond the unit-disk requirement and encode MWIS problems on arbitrarily connected graphs into two-dimensional atom arrays. We show the strategy of Ref. [19] in Fig. 1. In this scheme, each graph vertex is mapped into a vertex-wire: a one-dimensional array of nearest-neighbor-blockaded atoms, whose ground states, in the $\Omega/\Delta\rightarrow 0$ limit, are the $\mathbb{Z}_{2}$ ordered states [49]. The absence or the presence of an edge between two vertices is encoded by placing the crossing or the CWE gadget at the intersection of the corresponding vertex-wires, respectively [see Fig. 1(a)].
II.2.1 Single vertex-wire
@@
We consider a vertex-wire of $N$ atoms, with $N$ even [see Fig. 1(b)]. This is described by the Hamiltonian in Eq. (2), with homogeneous detuning $\Delta>0$ in the bulk and smaller detuning on the two boundary atoms: $\Delta_{1}=\Delta-\epsilon_{i}>0$ and $\Delta_{N}=\Delta-\epsilon_{f}>0$. In the blockade limit, the binary states ${\mu\in\{0,1\}}$ of the graph vertex correspond to the lowest two energy states of the Hamiltonian for $\Omega=0$, namely $|0\rangle_{l}=|0101\dots 01\rangle$ and $|1\rangle_{l}=|1010\dots 10\rangle$. We refer to these two $\mathbb{Z}_{2}$ ordered states as logical states. The corresponding energies are $-\Delta N/2+\epsilon_{f}$ and $-\Delta N/2+\epsilon_{i}$, respectively. The weight $\delta$ of the graph vertex (also referred to as wire weight) is mapped to the difference of these energies ${\delta=\epsilon_{f}-\epsilon_{i}}$.
II.2.2 Two intersecting vertex-wires
@@
The absence (presence) of an edge between two vertices is encoded by placing a crossing(-with-edge) gadget at the intersection between the corresponding vertex-wires. In reference to Fig. 1(c,d), we set the following terminology and notation. We use the label $w$ to indicate the vertex-wires: the horizontal one is denoted as $w=x$ and is oriented from left to right; the vertical one is denoted as $w=y$ and is oriented from top to bottom. The gadget divides each vertex-wire into two legs. The total number of atoms $N_{w}$ of the vertex-wire $w$ is the sum of the atoms of its two legs. We indicate the length of the first leg of the vertex-wire $w$ (the leg before the gadget, according to its orientation) with $m_{w}$. We choose both $N_{w}$ and $m_{w}$ to be even. We consider homogeneous detuning $\Delta>0$ on the whole system, except for $\Delta-\epsilon_{i,w}>0$ and $\Delta-\epsilon_{f,w}>0$ (with $w=x,y$) on the boundary atoms of the two vertex-wires, and $\Delta_{g}\geq\Delta$ on the gadget atoms [see Fig. 1(c,d)].
@@
The four logical states of the intersecting vertex-wires (i.e., the product of the logical states of the individual vertex-wires) are denoted with $|\mu_{x}\mu_{y}\rangle_{l}$, where ${\mu_{w}\in\{0,1\}}$ refers to the logical state of the $w$ vertex-wire. For $\Omega=0$ and equal boundary detuning (${\delta_{x}=\delta_{y}=0}$), these states constitute the ground states of the crossing setup [see Fig. 1(c)(right)], while the $|11\rangle_{l}$ state is prohibited in the CWE setup by the blockade mechanism [see Fig. 1(d)(right)]. Therefore, while the crossing gadget does not introduce any constraint among the logical degrees of freedom of the vertex-wires, the CWE gadget effectively imposes $\mu_{x}\mu_{y}=0$. These gadgets thus represent the absence and the presence of an edge between two vertices of the original graph, respectively.
@@ II.3.1 Standard protocol
The standard protocol, shown in Fig. 2, is commonly implemented in experiments [4]. All atoms are initialized in the $|0\rangle$ state, which is the ground state of the system for $\Delta(0)=-\Delta_{0}<0$ and $\Omega(0)=0$. Then, $\Omega(\tau)$ is ramped up to a finite value $\Omega_{0}$ and eventually down to zero, while $\Delta(\tau)$ is increased up to $\Delta_{0}$. The boundary detuning of each vertex-wire are varied proportionally to $\Delta(\tau)$, keeping their difference fixed to $(\epsilon_{f}(\tau)-\epsilon_{i}(\tau))/\Delta(\tau)=\delta/\Delta_{0}$, with $\delta$ being the wire weight.
@@ II.3.2 Logical protocol
As an alternative to the standard protocol, we consider the logical protocol shown in Fig. 2, which can be implemented with the same experimental resources as the standard one, and whose performance can be analyzed using perturbation theory. In this protocol, all vertex-wires are initialized in the logical state $|0\rangle_{l}$, which is the ground state of the system for $\Delta(0)=\Delta_{0}>0$, $\Omega(0)=0$, and ${\epsilon_{i}(0)>\epsilon_{f}(0)}$. Then, $\Omega(\tau)$ is ramped up to a finite value $\Omega_{0}$ and eventually down to zero, while $\Delta(\tau)$ is kept constant along the protocol. On each vertex-wire, $\epsilon_{i}(\tau)$ is ramped down while $\epsilon_{f}(\tau)$ is ramped up to satisfy the weight condition at final time: $\epsilon_{f}(T)-\epsilon_{i}(T)=\delta$, with $\delta$ being the wire weight.
@@ III Adiabatic performance with standard protocol
In this section, we investigate the adiabatic performance of two vertex-wires intersecting via a gadget when considering the standard protocol of Fig. 2. To perform this analysis, we develop an intersecting tensor network method (see details in Appendix A) and conduct a variational study using DMRG [64, 65, 66].
@@ III.1 Crossing setup
We consider the ground state of two vertex-wires intersecting via the crossing gadget [see Fig. 1(c)]. Depending on the value of $\Delta_{g}/\Delta$, the gadget may or may not host a Rydberg excitation. This effectively imposes different boundary constraints on the legs, leading to various ground-state regimes. The different regimes can be classified from the ground-state properties, whose changes can be tracked by computing the fidelity susceptibility [67, 68]:
$\chi=-\frac{2}{\delta\lambda^{2}}\ln\left|\langle\mathrm{GS}_{\lambda}|\mathrm{GS}_{\lambda+\delta_{\lambda}}\rangle\right|,$ (3)
@@
Figure 3: Analysis of the crossing setup. (a) Fidelity susceptibility $\chi$ in Eq. (3) (computed with ${\delta\lambda=0.05}$), as a function of $\Delta_{g}/\Omega$ and $\Delta/\Omega$. We consider two vertex-wires of $N=100$ atoms intersecting at the middle (${m_{x}=m_{y}=N/2}$). (b) Scaling of the minimum gap $\delta E_{m}$ (normalized to the maximum Rabi frequency $\Omega_{0}$) along the standard protocol in Fig. 2 as a function of $N$, for different values of ${\Delta_{g}/\Delta}$.
@@
As a dominant feature, we observe a high susceptibility around $\Delta/\Omega\sim 1.3$, where the system undergoes a phase transition between the disordered phase and the ordered one. This transition corresponds to the four legs undergoing (individually, or in pairs) the $\mathbb{Z}_{2}$ spontaneous symmetry breaking observed in one-dimensional Rydberg chains with nearest-neighbor blockade [49, 69, 70]. More precisely, for $\Delta_{g}<0$, the gadget is “empty”, i.e., all the gadget atoms are in the ground state. This has the effect of decoupling the four legs, which behave as independent Rydberg chains with open boundary conditions. In the transition from disorder to order, they thus individually break the $\mathbb{Z}_{2}$ symmetry, and the system enters an “inactive” ordered regime. Here, the ground state is given by one of the $2^{4}$ combinations of the $\mathbb{Z}_{2}$ states of the legs (and empty gadget). We describe the overall transition as a spontaneous breaking of a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. In the opposite limit of $\Delta_{g}\gtrsim\Delta$, the gadget always hosts a Rydberg excitation. This has the effect of connecting opposite legs, which behave as single vertex-wires. In this case, in the transition from disorder to order, the system spontaneously breaks a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry and enters an “active” ordered regime. Here, the ground state is given by one of the $2^{2}$ combinations of the $\mathbb{Z}_{2}$ states of the vertex-wires.
@@
We now consider the scaling of the minimum gap, located around the transition from the disordered phase to the active regime at $\Delta/\Omega\sim 1.3$, along the standard protocol. We target the $|11\rangle_{l}$ state, encoding the solution to the trivial MWIS problem for a graph of two disconnected vertices with equal weights, by setting ${\epsilon_{i}(\tau)=0.8\Delta(\tau)}$ and ${\epsilon_{f}(\tau)=0.2\Delta(\tau)}$ on both vertex-wires. In Fig. 3(b), we show that $\delta E_{m}$ scales as $1/L$ for $\Delta_{g}/\Delta>1$ (Ising transition, see Appendix D), and it still closes polynomially for $\Delta_{g}=\Delta$. We thus conclude that the crossing gadget does not constitute an obstruction to the adiabatic protocol. This result is shown for the symmetric configuration, but we confirmed that it is independent of where the two vertex-wires intersect.

### v2

@@ Abstract
Quantum adiabatic optimization seeks to solve combinatorial problems using quantum dynamics, requiring the Hamiltonian of the system to align with the problem of interest. However, these Hamiltonians are often incompatible with the native constraints of quantum hardware, necessitating encoding strategies to map the original problem into a hardware-conformant form. While the classical overhead associated with such mappings is easily quantifiable and typically polynomial in problem size, it is much harder to quantify their overhead on the quantum algorithm, e.g., the transformation of the adiabatic timescale. In this work, we address this challenge on the concrete example of the encoding scheme proposed in [Nguyen et al., PRX Quantum 4, 010316 (2023)], which is designed to map optimization problems on arbitrarily connected graphs into Rydberg atom arrays. We consider the fundamental building blocks underlying this encoding scheme and determine the scaling of the minimum gap with system size along adiabatic protocols. Even when the original problem is trivially solvable, we find that the encoded problem can exhibit an exponentially closing minimum gap. We show that this originates from a quantum coherent effect, which gives rise to an unfavorable localization of the ground-state wavefunction. On the QuEra Aquila neutral atom machine, we observe such localization and its effect on the success probability of finding the correct solution to the encoded optimization problem. Finally, we propose quantum-aware modifications of the encoding scheme that avoid this quantum bottleneck and lead to an exponential improvement in the adiabatic performance. This highlights the crucial importance of accounting for quantum effects when designing strategies to encode classical problems onto quantum platforms.
@@ I Introduction
The past few decades have seen a growing interest in harnessing the dynamics of many-body quantum systems to explore potential quantum advantage in solving combinatorial optimization problems [1, 2, 3, 4]. For problems that can be reformulated in terms of a classical spin Hamiltonian whose ground state represents the solution, quantum adiabatic optimization stands out as one of the most promising quantum algorithms for pursuing such an advantage [5, 6, 7, 8, 9, 10, 11]. Here, the Hamiltonian of the system is adiabatically evolved from one with a known and easy-to-prepare ground state to the target one. In this way, the system follows the instantaneous ground state of the time-dependent Hamiltonian throughout the evolution and finally reaches the solution to the optimization problem.
In practice, many problems of interest do not naturally conform with the native constraints of the quantum hardware. Therefore, additional encoding strategies to map them into a hardware-conformant form are required [12, 13]. While the classical overhead associated with such strategies is easily quantifiable and typically polynomial in problem size, understanding their impact on the quantum dynamics of the adiabatic algorithm remains an outstanding challenge.
@@
In this work we address this question, focusing on the example of quantum optimizers based on Rydberg atom arrays [14, 4, 15, 16, 17, 18, 19]. Due to the blockade mechanism, Rydberg atom arrays naturally encode the paradigmatic maximum independent set problem on a specific class of graphs, so-called unit-disk graphs [14], and experiments on these graphs have already shown promising results [4]. Several encoding strategies to go beyond the native hardware restriction of unit-disk graphs have been proposed in the literature  [20, 21, 22, 23, 24, 25, 26]. In particular, Ref. [21] proposed an encoding strategy that maps an optimization problem on an arbitrarily connected graph into an equivalent maximum weighted independent set (MWIS) problem on a unit-disk graph, which is directly implementable with Rydberg atom arrays (see Fig. 1). Here, the binary degree of freedom associated with each graph vertex is encoded in the $\mathbb{Z}_{2}$-ordered states of a vertex wire of Rydberg atoms. The absence (presence) of an edge between a pair of vertices is reproduced by placing a crossing (crossing-with-edge) gadget at the intersection between the corresponding vertex wires. So far, this strategy has been mainly investigated on the classical level, and the quantum adiabatic performance, in particular for large systems, remains an outstanding open question.
@@
In this work, we investigate the quantum adiabatic performance of the main building blocks of this encoding scheme, consisting of two vertex wires intersecting either via the crossing or the crossing-with-edge (CWE) gadget, as depicted in Fig. 1(c,d). When the two vertex wires intersect via the crossing-with-edge gadget, we find that an exponential closing of the minimum gap can occur for certain geometrical configurations, leading to exponentially long algorithm runtimes for large systems. We show that this adversarial gap closing originates from unfavorable localization of the ground-state wavefunction along the adiabatic protocol. Guided by these insights, we propose alternative strategies to overcome such a localization and ensure a polynomial closing of the minimum gap, thereby exponentially improving the adiabatic performance of the encoding scheme. In particular, we show how localization can be avoided by (i) systematically varying the length of the intersecting vertex wires, or (ii) modifying the structure of the crossing-with-edge gadget.
@@
This work is organized as follows. In Sec. II, we revisit the encoding of the MWIS problem on arbitrarily connected graphs into Rydberg atom arrays, as proposed in Ref. [21]. After detailing its building blocks (namely, the crossing and the CWE setup), we introduce the two adiabatic protocols that we consider in this work, referred to as standard and logical protocols. In Sec. III, we employ the density-matrix renormalization group (DMRG) method to quantify the adiabatic performance of the two setups along the standard protocol. In Sec. IV, we develop an effective theory to characterize their performance along the logical protocol and explain the origin of the wavefunction localization. In Sec. V, using the QuEra Aquila neutral atom machine [27], we observe such localization phenomenon and its effect on encoded MWIS problems. Finally, in Sec. VI, we present the strategies to overcome the localization of the ground-state wavefunction and achieve an exponential improvement in the adiabatic performance.
@@ II Rydberg encoding of maximum weighted independent set problems
Figure 1: Encoding scheme of the maximum weighted independent set problem on arbitrarily connected graphs into Rydberg atom arrays [21]. (a) Example for a non-unit-disk graph of five vertices $G=(V,E,\delta)$. (b) A vertex wire is a one-dimensional array of nearest-neighbor-blockaded atoms. (Left) The atoms are labeled from $1$ to $N$, with $N$ even. The detuning in the bulk is denoted by $\Delta$; the detuning on the initial and final atom by $\Delta-\epsilon_{i}$ and $\Delta-\epsilon_{f}$. The weight of a graph vertex $\delta$ is mapped to the boundary detuning difference: $\delta=\epsilon_{f}-\epsilon_{i}$. (Right) The logical states encoding the binary degrees of freedom of a graph vertex are the $\mathbb{Z}_{2}$-ordered states of the vertex wire. Here, the atoms in the ground (Rydberg) state are depicted in white (black). The (c) crossing and (d) CWE setups consist of two vertex wires intersecting via the crossing gadget (the four atoms in gray) and the CWE gadget (the three atoms in gray, labeled as in the inset). (Left) The atoms along the horizontal ($w=x$) and vertical ($w=y$) vertex wire are labeled from $1$ to $N_{w}$; the position of the last atom before the gadget is denoted by $m_{w}$. $N_{w}$ and $m_{w}$ are chosen to be even. The bulk detuning is denoted by $\Delta$; the detuning on the initial and the final atom of the $w$ vertex wire by $\Delta-\epsilon_{i,w}$ and $\Delta-\epsilon_{f,w}$; the detuning on the gadget atoms by $\Delta_{g}$. (Right) The logical states of the crossing setup are the four possible combinations of logical states of the vertex wires. In the CWE setup, the logical state $|11\rangle_{l}$ is forbidden due to the blockade mechanism.
@@ II.2 MWIS encoding for arbitrarily connected graphs
A quadratic overhead in the atom number makes it possible to go beyond the unit-disk requirement and encode MWIS problems on arbitrarily connected graphs into two-dimensional atom arrays. We show the strategy of Ref. [21] in Fig. 1. In this scheme, each graph vertex is mapped into a vertex wire: a one-dimensional array of nearest-neighbor-blockaded atoms, whose ground states, in the $\Omega/\Delta\rightarrow 0$ limit, are the $\mathbb{Z}_{2}$-ordered states [51]. The absence or the presence of an edge between two vertices is encoded by placing the crossing or the CWE gadget at the intersection of the corresponding vertex wires, respectively [see Fig. 1(a)].
II.2.1 Single vertex wire
@@
We consider a vertex wire of $N$ atoms, with $N$ even [see Fig. 1(b)]. This is described by the Hamiltonian in Eq. (2), with homogeneous detuning $\Delta>0$ in the bulk and smaller detuning on the two boundary atoms: $\Delta_{1}=\Delta-\epsilon_{i}>0$ and $\Delta_{N}=\Delta-\epsilon_{f}>0$. In the blockade limit, the binary states ${\mu\in\{0,1\}}$ of the graph vertex correspond to the lowest two energy states of the Hamiltonian for $\Omega=0$, namely $|0\rangle_{l}=|0101\dots 01\rangle$ and $|1\rangle_{l}=|1010\dots 10\rangle$. We refer to these two $\mathbb{Z}_{2}$-ordered states as logical states. The corresponding energies are $-\Delta N/2+\epsilon_{f}$ and $-\Delta N/2+\epsilon_{i}$, respectively. The weight $\delta$ of the graph vertex (also referred to as wire weight) is mapped to the difference of these energies ${\delta=\epsilon_{f}-\epsilon_{i}}$.
II.2.2 Two intersecting vertex wires
@@
The absence (presence) of an edge between two vertices is encoded by placing a crossing(-with-edge) gadget at the intersection between the corresponding vertex wires. In reference to Fig. 1(c,d), we set the following terminology and notation. We use the label $w$ to indicate the vertex wires: the horizontal one is denoted as $w=x$ and is oriented from left to right; the vertical one is denoted as $w=y$ and is oriented from top to bottom. The gadget divides each vertex wire into two legs. The total number of atoms $N_{w}$ of the vertex wire $w$ is the sum of the atoms of its two legs. We indicate the length of the first leg of the vertex wire $w$ (the leg before the gadget, according to its orientation) with $m_{w}$. We choose both $N_{w}$ and $m_{w}$ to be even. We consider homogeneous detuning $\Delta>0$ on the whole system, except for $\Delta-\epsilon_{i,w}>0$ and $\Delta-\epsilon_{f,w}>0$ (with $w=x,y$) on the boundary atoms of the two vertex wires, and $\Delta_{g}\geq\Delta$ on the gadget atoms [see Fig. 1(c,d)].
@@
The four logical states of the intersecting vertex wires (i.e., the product of the logical states of the individual vertex wires) are denoted with $|\mu_{x}\mu_{y}\rangle_{l}$, where ${\mu_{w}\in\{0,1\}}$ refers to the logical state of the $w$ vertex wire. For $\Omega=0$ and equal boundary detuning (${\delta_{x}=\delta_{y}=0}$), these states constitute the ground states of the crossing setup [see Fig. 1(c)(right)], while the $|11\rangle_{l}$ state is prohibited in the CWE setup by the blockade mechanism [see Fig. 1(d)(right)]. Therefore, while the crossing gadget does not introduce any constraint among the logical degrees of freedom of the vertex wires, the CWE gadget effectively imposes $\mu_{x}\mu_{y}=0$. These gadgets thus represent the absence and the presence of an edge between two vertices of the original graph, respectively.
@@ II.3.1 Standard protocol
The standard protocol, shown in Fig. 2, is commonly implemented in experiments [4]. All atoms are initialized in the $|0\rangle$ state, which is the ground state of the system for $\Delta(0)=-\Delta_{0}<0$ and $\Omega(0)=0$. Then, $\Omega(\tau)$ is ramped up to a finite value $\Omega_{0}$ and eventually down to zero, while $\Delta(\tau)$ is increased up to $\Delta_{0}$. The boundary detuning of each vertex wire are varied proportionally to $\Delta(\tau)$, keeping their difference fixed to $(\epsilon_{f}(\tau)-\epsilon_{i}(\tau))/\Delta(\tau)=\delta/\Delta_{0}$, with $\delta$ being the wire weight.
@@ II.3.2 Logical protocol
As an alternative to the standard protocol, we consider the logical protocol shown in Fig. 2, which can be implemented with the same experimental resources as the standard one, and whose performance can be analyzed using perturbation theory. In this protocol, all vertex wires are initialized in the logical state $|0\rangle_{l}$, which is the ground state of the system for $\Delta(0)=\Delta_{0}>0$, $\Omega(0)=0$, and ${\epsilon_{i}(0)>\epsilon_{f}(0)}$. Then, $\Omega(\tau)$ is ramped up to a finite value $\Omega_{0}$ and eventually down to zero, while $\Delta(\tau)$ is kept constant along the protocol. On each vertex wire, $\epsilon_{i}(\tau)$ is ramped down while $\epsilon_{f}(\tau)$ is ramped up to satisfy the weight condition at final time: $\epsilon_{f}(T)-\epsilon_{i}(T)=\delta$, with $\delta$ being the wire weight.
@@ III Adiabatic performance with standard protocol
In this section, we investigate the adiabatic performance of two vertex wires intersecting via a gadget when considering the standard protocol of Fig. 2. To perform this analysis, we develop an intersecting tensor-network method (see details in Appendix A) and conduct a variational study using DMRG [66, 67, 68].
@@ III.1 Crossing setup
We consider the ground state of two vertex wires intersecting via the crossing gadget [see Fig. 1(c)]. Depending on the value of $\Delta_{g}/\Delta$, the gadget may or may not host a Rydberg excitation. This effectively imposes different boundary constraints on the legs, leading to various ground-state regimes. The different regimes can be classified from the ground-state properties, whose changes can be tracked by computing the fidelity susceptibility [69, 70]:
$\chi=-\frac{2}{\delta\lambda^{2}}\ln\left|\langle\mathrm{GS}_{\lambda}|\mathrm{GS}_{\lambda+\delta{\lambda}}\rangle\right|,$ (3)
@@
Figure 3: Analysis of the crossing setup. (a) Fidelity susceptibility $\chi$ in Eq. (3) (computed with ${\delta\lambda=0.05}$), as a function of $\Delta_{g}/\Omega$ and $\Delta/\Omega$. We consider two vertex wires of $N=100$ atoms intersecting at the middle (${m_{x}=m_{y}=N/2}$). (b) Scaling of the minimum gap $\delta E_{m}$ (normalized to the maximum Rabi frequency $\Omega_{0}$) along the standard protocol in Fig. 2 as a function of $N$, for different values of ${\Delta_{g}/\Delta}$.
@@
As a dominant feature, we observe a high susceptibility around $\Delta/\Omega\sim 1.3$, where the system undergoes a phase transition between the disordered phase and the ordered one. This transition corresponds to the four legs undergoing (individually, or in pairs) the $\mathbb{Z}_{2}$ spontaneous symmetry breaking observed in one-dimensional Rydberg chains with nearest-neighbor blockade [51, 71, 72]. More precisely, for $\Delta_{g}<0$, the gadget is “empty”, i.e., all the gadget atoms are in the ground state. This has the effect of decoupling the four legs, which behave as independent Rydberg chains with open boundary conditions. In the transition from disorder to order, they thus individually break the $\mathbb{Z}_{2}$ symmetry, and the system enters an “inactive” ordered regime. Here, the ground state is given by one of the $2^{4}$ combinations of the $\mathbb{Z}_{2}$ states of the legs (and empty gadget). We describe the overall transition as a spontaneous breaking of a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. In the opposite limit of $\Delta_{g}\gtrsim\Delta$, the gadget always hosts a Rydberg excitation. This has the effect of connecting opposite legs, which behave as single vertex wires. In this case, in the transition from disorder to order, the system spontaneously breaks a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry and enters an “active” ordered regime. Here, the ground state is given by one of the $2^{2}$ combinations of the $\mathbb{Z}_{2}$ states of the vertex wires.
@@
We now consider the scaling of the minimum gap, located around the transition from the disordered phase to the active regime at $\Delta/\Omega\sim 1.3$, along the standard protocol. We target the $|11\rangle_{l}$ state, encoding the solution to the trivial MWIS problem for a graph of two disconnected vertices with equal weights, by setting ${\epsilon_{i}(\tau)=0.8\Delta(\tau)}$ and ${\epsilon_{f}(\tau)=0.2\Delta(\tau)}$ on both vertex wires. In Fig. 3(b), we show that $\delta E_{m}$ scales as $1/L$ for $\Delta_{g}/\Delta>1$ (Ising transition, see Appendix D), and it still closes polynomially for $\Delta_{g}=\Delta$. We thus conclude that the crossing gadget does not constitute an obstruction to the adiabatic protocol. This result is shown for the symmetric configuration, but we confirmed that it is independent of where the two vertex wires intersect.